a major computer company has just purchased the new operating system you created you have so much money that you decide to give $2,000,000 away. If you give $50 away every hour , how long will it take you in years ?
1
Expert's answer
2011-11-10T09:48:57-0500
It will take H = 2,000,000 / 50 = 40,000 hours to to give $2,000,000 away.
There are 24*365 = 8760 hours in a year (24 hours in a day and 365 days in a year).
So, it will take Y = H/8760 = 40,000/8760 ≈ 4.566 years.